# Efficient Online Reinforcement Learning Fine-Tuning Need Not Retain Offline Data

### v2

@@ 5.1 Baselines and Experimental Setup
Figure 8: KL divergence of the pre-trained policy and Q-function with fine-tuned ones show that WSRL did not forget its pre-training. The four plots show $D_{\mathrm{KL}}(\pi_{\psi}^{\mathrm{pre}}||\pi_{\psi})$ and $D_{\mathrm{KL}}(\mathrm{softmax}(Q_{\theta}^{\mathrm{pre}})||\mathrm{softmax}(\pi_{\psi}))$ on online and offline state distributions. This plot aggregates the KL divergence of six different environments shown separately in Figure 19 and 20, along with a more detailed discussion in Appendix E.
@@ 5.4 How critical is the warmup phase?
We find that the warmup phase is crucial for fine-tuning with online RL. As shown in Figure 10, WSRL without warmup performs significantly worse in three different environments. Moreover, we find that using such simple warmup scheme is better than WSRL initializing with the same number of transitions from the offline dataset or retaining the offline dataset all together (See Appendix J and K). We hypothesize that warmup helps because it helps mitigate the distribution shift and prevents divergence and the “downward spiral” at the beginning of fine-tuning. As shown in Figure 11 shows, learned Q-values do not diverge to overly pessimistic values when warmup is employed and the TD losses remain small on the offline data as well, avoiding the “downward spiral” in Section 3. See more detailed discussion in Appendix G.
@@ Appendix I Implementation Details
Appendix J Warmup with Transitions from the Offline Dataset
@@
Appendix K WSRL with Offline Data Retention
@@
Appendix L Ablating the Effects of Q-ensemble and Layer Normalization
@@
Appendix M WSRL with Different Update-to-Data Ratios
@@
Appendix N WSRL with Varying Levels of Offline Policy

### v3

@@ 5.1 Baselines and Experimental Setup
Figure 8: KL divergence of the pre-trained policy and Q-function with fine-tuned ones show that WSRL did not forget its pre-training. The four plots show $D_{\mathrm{KL}}(\pi_{\psi}^{\mathrm{pre}}||\pi_{\psi})$ and $D_{\mathrm{KL}}(\mathrm{softmax}(Q_{\theta}^{\mathrm{pre}})||\mathrm{softmax}(\pi_{\psi}))$ on online and offline state distributions. This plot aggregates the KL divergence of six different environments shown separately in Figure 21 and 22, along with a more detailed discussion in Appendix E.
@@ 5.4 How critical is the warmup phase?
We find that the warmup phase is crucial for fine-tuning with online RL. As shown in Figure 10, WSRL without warmup performs significantly worse in three different environments. Moreover, we find that using such simple warmup scheme is better than WSRL initializing with the same number of transitions from the offline dataset or retaining the offline dataset all together (See Appendix K and L). We hypothesize that warmup helps because it helps mitigate the distribution shift and prevents divergence and the “downward spiral” at the beginning of fine-tuning. As shown in Figure 11 shows, learned Q-values do not diverge to overly pessimistic values when warmup is employed and the TD losses remain small on the offline data as well, avoiding the “downward spiral” in Section 3. See more detailed discussion in Appendix G.
@@ 5.6 How important is it to initialize the policy, value function, and both?
5.7 How does WSRL perform on Real World Robotic Tasks?
Task WSRL SERL Offline Pretrained
@ 18 min @ 50 min Cal-QL
Franka Peg Insertion 20/20 $0/20$ $13/20$
Table 1: WSRL enables efficient real-world RL fine-tuning. WSRL quickly improves upon offline RL pre-training and fine-tunes much faster than RLPD/SERL. WSRL takes 11 min to collect 5k steps of warmup data, before only doing 7 min of online RL to master the task. Evaluated with 20 different initial poses.
Real-world robot and task setup. We use a 7 DoF Franka Emika Panda arm with end-effector control at 10 Hz. As shown in Figure 14, we use two wrist cameras images and robot proprioceptive state as input. The task, “Franka Peg Insertion”, requires rotating, navigating, and inserting a plastic block with pegs into the corresponding holes in a box-shaped frame, with 3D printable objects from Luo et al. [33].
Figure 14: Franka robot setup for peg insertion experiment.
Data collection and environment. Since there is no readily available pre-training dataset for the peg insertion task, we collect a small dataset of around $17$K transitions. This dataset is collected by saving the replay buffer during an online RL run. As a result, our dataset contains a mixed quality data that typically constitute expert-replay datasets [12]. We choose to collect such mixed quality data over tele-operated human demonstration data to test the offline RL algorithm’s ability to learn from suboptimal data. More details in Appendix J.
Real-world results. We compare WSRL to SERL [31], a strong prior approach that incorporates offline data into efficient online RL training on real robots, building on the RLPD [4] algorithm. We train both methods online for $20$K environment steps. WSRL is initialized with an offline checkpoint trained by CalQL on our offline dataset. For an apples-to-apples comparison, we initialize SERL’s replay buffer with the same offline dataset. We evaluate on 20 trials starting from initial states evenly distributed around the goal pose, with results in Table 1.
We find that in just 18 minutes of real-time RL training (roughly $5$K warmup + $3$K actor steps, 7 minutes without including warmup), WSRL is able to fine-tune from an offline performance of 13/20 to a perfect 20/20. In comparison, we find that SERL/RLPD achieves 0/20 at the end of  $50$ minutes ($20$k actor step). Qualitatively, we observe that SERL often gets stuck in the out-of-distribution positions, simply stopping and failing to progress further (See Figure 15 for an example). Even when operating within the support of the offline distribution, SERL frequently hovers over various positions without committing to an insertion path, showing that it was unable to learn from the offline dataset. In comparison, WSRL is able to effectively utilize its offline RL pre-training and bootstrap from that knowledge to quickly fine-tune online.
WSRL Successful Insertion
Starting position successfully navigates and inserts
RLPD Failure Mode
Starting position fails to insert the peg
Figure 15: WSRL (in green) learns an expert-level policy and can insert successfully from any location. SERL (in red) fails to learn a successful policy from the offline dataset and drifts into out-of-distribution positions, where it then stalls.
@@ Appendix I Implementation Details
Appendix J Implementation Details for Franka Robot
Robot Setup. We run our experiments on the Franka robot arm using the SERL software suite [31]. During online RL fine-tuning, SERL runs data collection and agent update as asynchronous processes. For the peg insertion task, we fix the robot gripper to be always closed. Following  Luo et al. [31, 32], we apply restrict the robot’s end effector positions to be within a certain bounding box as described below. Our policy uses a pre-trained resnet encoder and optimizes a sparse reward provided by a binary classifer pre-trained on manually-collected 200 positive and 500 negative examples.
Data Collection. We record the replay buffer when running SERL initialized with $20$ pre-collected expert demos, and use human interventions to guide and speed up learning during online RL [32] for fast data collection (When human intervention is not used, we find that SERL takes over an hour to solve the peg insertion task, consistent with that reported by  Luo et al. [31]).
Environment Details. Following  Luo et al. [31, 32], we restrict the robot’s end effector positions to be within a certain area centered at the insertion position. During data collection for the offline dataset, we loosen the restriction and use a bigger bounding box that spans the perimeter of the green box to collect diverse data. We use a discount of $0.98$, and uses random resets within the bounding box for online finetuning.
Hyperparameters. We use 5K warmup steps and follow the SERL [31] SAC implementation. For RLPD, CalQL, and WSRL, we use a 2-block resnet MLP, layer-norm, and an critic-to-actor ratio of $4$ with batch size $256$, and a learning rate of $43e-$ 4. We use an ensemble of $10$ Q-functions and predict the Q value by randomly sub-sample $2$ and taking the min. For offline CalQL, we use a fixed CQL regularizer weight $\alpha=5$, and train for 40k gradient steps. Code for robot experiments can be found at https://github.com/zhouzypaul/wsrl-robot.
Appendix K Warmup with Transitions from the Offline Dataset
@@
Appendix L WSRL with Offline Data Retention
@@
Appendix M Ablating the Effects of Q-ensemble and Layer Normalization
@@
Appendix N WSRL with Different Update-to-Data Ratios
@@
Appendix O WSRL with Varying Levels of Offline Policy